Just solve A and B please!

Velocity addition A train with proper length L' t
Answer #1

a) First, we have to find the conductor speed wheen he is moving trough the train, in the trains frame, which mean that we are going to find that speed beeing inside the train with him. This is:

V_{c}^{'} = \frac{L^{'}}{t_{c}^{'}}

Now, lets consider that we're off the train. We see the train traveling at a speed equals Vt, and we want to determine the speed of the conductor when he is walking from front to back and from back to front. If in an initial moment when we see the train drive away from us, we see the conductor walking from:

  • From back to front: then his speed respect us (observer O) is equals: V_{cO} = \frac{L^{'}}{t_{c}^{'}} + Vt
  • From front to back: then his speed respect us (observer O) is equals: V_{cO} = Vt - \frac{L^{'}}{t_{c}^{'}}

b) Now imagine that the conductor kicked all that things out of the train. This makes objects describe a parabolic path relative to the floor of the train at different angles as we know. If we see as the observer O (out of the train) all those objects flying, we will see than they described another parabolic path because of the train lineal speed (Vt). We can say then, that we have an horizontal speed component and a vertical speed component, which is used to find "w" (speed relative O's frame) with the Pythagoras theorem:

W = \sqrt{Vt^{2} + S_{o}^{'}^{2}}

Finally, we can also say that, respect an observer O, the horizontal speed component is Vt (adyacent side) and the vertical speed component is So (opossite side). by using trigonometry:

\tan \theta _{i} = \frac{S_{o}^{'}}{Vt} \rightarrow \arctan \frac{S_{o}^{'}}{Vt} = \theta _{i}